Assuming that the planets are moving in circular orbits, apply Kepler's laws to show that the acceleration of a planet is inversely proportional to the square of its distance from the sun.
1
Expert's answer
2021-02-15T17:43:46-0500

Answer

Let us suppose

the planets are moving in circular orbits.

So gravitational force can be written as

"F=\frac{GMm}{r^2}" ...... 1

Force is given as

"F=ma" ...... 2

Using both equation

"F=ma=\frac{GMm}{r^2}"

"a=\frac{GM}{r^2}"

So acceleration

"a\propto\frac{1}{r^2}"